OpenStudy (anonymous):

Find the y-intercept in this linear equation: 2y - 4x = -12

OpenStudy (anonymous):

change the equation to y=mx+b form then b will be y intercept

OpenStudy (anonymous):

add 4x to both sides then devide by 2

OpenStudy (anonymous):

2y = -12 + 4x 2y/2 = -12/2 +4x/2 y = -6 + 2x I'm so confused

OpenStudy (anonymous):

or would y = -6?

OpenStudy (zzr0ck3r):

correct

OpenStudy (radar):

Just let y=0 and you see that -4x=-12, so the x intercept is 3 let x=0 and y=-6

OpenStudy (anonymous):

So, if they don't give you a "x" it is automatically set to 0?

OpenStudy (radar):

If no x is given, then it is a horizontal line where y is a constant.

OpenStudy (anonymous):

If i'm finding the x-intercept it would be: 2y - 4x = -12 2(0) -4x = -12 -4x/-4 = -12/-4 x = 3 ?

OpenStudy (radar):

Good.

OpenStudy (anonymous):

Thank you very much, all of you.

OpenStudy (radar):

X intercept, is the point on a line that intercepts the x axis, you let y=0 to find the x intercept, and let x=0 to find the y intercept.